(√(c-5)2 + (c2-2))= 23


1
Expert's answer
2021-07-28T16:22:15-0400

"\sqrt{(c-5)^2+(c^2-2)}=23"

"(c-5)^2+(c^2-2)=23^2"

"2c^2-10c-506=0"

"c^2-5c-253=0"

"c_1=\frac{5+\sqrt{1037}}{2}"

"c_2=\frac{5-\sqrt{1037}}{2}"